OpenStudy (anonymous):

the square of x is = to 4 times the square of y. If x is 1 more than twice y , what is the value of x?

OpenStudy (anonymous):

(2y + 1)^2 = 4y^2

OpenStudy (anonymous):

2y + 1 = 4y^2

OpenStudy (anonymous):

2y + 1 = 4y*

OpenStudy (blueraverwolf):

X=1/2

OpenStudy (anonymous):

can i do that

OpenStudy (anonymous):

report raver plz tired of trolls

OpenStudy (blueraverwolf):

what are you talking about

OpenStudy (anonymous):

x = 4y^2 x = 2y +1 4y^2 = 2y +1 4y^2 - 2y -1 = 0 quadratic equation to solve for y

OpenStudy (anonymous):

u cant help by giving out answers ,... kid just get out .. obv common sense . OBV TROLL OBV

OpenStudy (anonymous):

whats wrong with my way?

OpenStudy (anonymous):

you read the question wrong it says that x = 2y +1 not x = (2y+1)^2

OpenStudy (anonymous):

was plugign it in the original equation

OpenStudy (anonymous):

x^2 = 4*y^2

OpenStudy (anonymous):

then subsitutting x

OpenStudy (anonymous):

with y

OpenStudy (anonymous):

ohh is says the square of x at the beginning, okay i didn't see that makes it much easier

OpenStudy (anonymous):

2y +1 = 4y 1 = 2y y = 1/2 plug that in for x = 2(1/2) +1 x = 2

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

the thing is thats not a answer

OpenStudy (anonymous):

the answer is 1/2 but i still don't get how anyone can get that

OpenStudy (anonymous):

sorry 1/4

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

x^2 = 4*y^2 x = 1 + 2y Substitute the second equation in the first (1+2y)^2 = 4y^2 1 + 4y + 4y^2 = 4y^2 4y = -1 y = -1/4 Substitute in y = -1/4 into the second equation: x = 1 - 2/4 x = 1/2

OpenStudy (anonymous):

can you tell me why my way is wrong plz